In 2018, Lew, who is single, earned wages of $386,724. a. What is the amount of Social Security tax withheld from his salary? b. What is the amount of Medicare tax (including additional 0.9% tax if any) withheld from his salary?
Answer #1
a
Social Security tax withheld 7960.80 =128400*6.2%
b
Medicare tax 5607.50 =386724*1.45%
Additional Medicare tax 1680.52 =(386724-200000)*0.9%
Amount of Medicare tax withheld 7288.02 or 7288.01
Note:
The 2018 wage limit for Social Security is $128400
Additional Medicare tax is payable for wages above $200000
Amounts have been rounded off to 2 decimal places